# Blackjack consecutive wins or losses

**URL:** <https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683>\
**Category:** Factual Questions\
**Created:** [June 29, 2004, 7:50am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683 "2004-06-29T07:50:13Z")\
**Posts on this page:** 10\
**Page:** 2

<div class="post-metadata">

**Author:** ![I\_Love\_Me\_Vol.I](https://sea3.discourse-cdn.com/straightdope/user_avatar/boards.straightdope.com/i_love_me_vol.i/32/521_2.png) [@I\_Love\_Me\_Vol.I](https://boards.straightdope.com/u/I_Love_Me_Vol.I)\
**Post date:** [June 29, 2004, 11:02pm UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/21 "2004-06-29T23:02:26Z")

</div>

> [@Josh\_dePlume](#):
>
> That’s not been my experience. I used to play single deck in Nevada all the time. The only unfavorable rule was dealer hits soft 17, which is about a 0.2% disadvantage to the player. That pales in comparison to the disadvantage of a large number of decks, which is about 0.6% (more decks is bad even for the non-counter; it changes things because you’re drawing without replacement).

When did you play? Years ago you could get single-deck with all the goodies. Not any more (except as \*\*Sam \*\* mentioned as a loss-leader promotion with $10 or 20 max bets–and even then VERY rarely).

In the games you played did you get to surrender? Did you get to double after splitting? Resplit aces? Split all cards up to 4 times? Double any 2 cards? Get a 3-2 BJ payoff?

I can get all those rules PLUS the dealer staying on soft 17 in Vegas tonight in a six-deck shoe game. That puts the casino edge at about .25 or so. Not the .6 you mention.

If you know of a current single deck game with ALL of the rules I mentioned above (even with the H17 you mentioned) for normal stakes (say up to $500 max) Please tell me where it is!

---

<div class="post-metadata">

**Author:** ![Little\_Nemo](https://sea3.discourse-cdn.com/straightdope/user_avatar/boards.straightdope.com/little_nemo/32/3120_2.png) [@Little\_Nemo](https://boards.straightdope.com/u/Little_Nemo)\
**Post date:** [June 30, 2004, 2:25am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/22 "2004-06-30T02:25:09Z")

</div>

> [@](#):
>
> So you disagree with what Uncommon Sense and I wrote?

Not completely. But for all practical purposes, yes. I’ll concede that losing three hands in a row might indicate the deck was now “hot”. (Although if the dealer had lost three hands in a row and you won them, wouldn’t the deck be equally “hot”?) But even perfect card counting only moves the odds a few percentage points in a player’s favor. Half-ass card counting like the OP describes is unlikely to move the odds past the built in house advantage.

---

<div class="post-metadata">

**Author:** ![Josh\_dePlume](https://avatars.discourse-cdn.com/v4/letter/j/bbe5ce/32.png) [@Josh\_dePlume](https://boards.straightdope.com/u/Josh_dePlume)\
**Post date:** [June 30, 2004, 3:06am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/23 "2004-06-30T03:06:40Z")

</div>

> [@I Love Me](#):
>
> In the games you played did you get to surrender? Did you get to double after splitting? Resplit aces? Split all cards up to 4 times? Double any 2 cards? Get a 3-2 BJ payoff?
> 
> I can get all those rules PLUS the dealer staying on soft 17 in Vegas tonight in a six-deck shoe game. That puts the casino edge at about .25 or so. Not the .6 you mention.

I played these single-deck games until about two years ago; I suppose it’s possible they’ve changed since then. I played mostly in a little town called Wendover on the Utah border, but found much the same conditions in Vegas, Reno, and some smaller places like Elko and Mesquite. Cards could be resplit indefinitely except for aces (resplitting aces is worth only 0.03% with a single deck). 3-2 BJ payoff (I’ve never seen anything else in practice, only in tables in books). Double on any 2 cards. No double after split, no surrender. So according to the tables I’ve looked at, e.g., in Griffin’s book and Thorp’s, that gives a house edge of about 0.2%.

Six decks is about a 0.6% disadvantage. Resplit of aces is about a 0.08% advantage with many decks, and double after split is 0.14%. The value of “conventional” surrender (as opposed to “early” surrender; I presume you’re not fortunate enough to have that) is about 0.085% under the conditions you describe (I’m getting these numbers from Griffin). So I get 0.295% house edge for your game.

---

<div class="post-metadata">

**Author:** ![I\_Love\_Me\_Vol.I](https://sea3.discourse-cdn.com/straightdope/user_avatar/boards.straightdope.com/i_love_me_vol.i/32/521_2.png) [@I\_Love\_Me\_Vol.I](https://boards.straightdope.com/u/I_Love_Me_Vol.I)\
**Post date:** [June 30, 2004, 5:08am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/24 "2004-06-30T05:08:30Z")

</div>

Josh–

Heck, if I could find a single deck game with a .20 edge for the house I’d love to play it. The only reason I play from a shoe is because those are the games with the best advantage I can find. I’d gladly sacrifice DAS and move to single deck if I knew where the game was. Especially if the penetration was better than half a deck.

---

<div class="post-metadata">

**Author:** ![EllisDee](https://sea3.discourse-cdn.com/straightdope/user_avatar/boards.straightdope.com/ellisdee/32/4531_2.png) [@EllisDee](https://boards.straightdope.com/u/EllisDee)\
**Post date:** [June 30, 2004, 5:18am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/25 "2004-06-30T05:18:26Z")

</div>

First, let me state clearly that I know nothing about blackjack or card counting. If your system is predicated on dependent trials, then my logic may not apply. (Though, I suspect, your trials are more independent than you think.)

> [@ccwaterback](#):
>
> I play a regular game until I lose 3 in a row. After losing 3 in a row, the odds are definitely in my favor. I then double my bet. Then win or lose, I am going to double my bet again, because the odds should still be in my favor **(odds of losing 4 out of 5 or 5 out of 5)**. Then win or lose, I’m going to back down to $10 and stay at $10 until I lose, then go back to where I started at $5. It just seems to me I am increasing my bet when the odds are in my favor, which should be a good thing.

The bolded (by me) part is irrelevant. It also happens to be the core of the gambler’s fallacy.

Let’s assume 50-50 wagers. Say, a coin flip.

What’s the chance to lose 3 in a row? 12.5%

So I lose 2 in a row. Now I should bump, because there’s a 100 - 12.5 = 87.5% chance to win, right?

Wrong. You just lost 2 in a row, which in itself only has a 25% chance to occur. That means that **you already realized** the low percentage chance you are trying to capitalize on. As in, you missed your chance.

Common sense tells you that losing 3 in a row only has a 12.5% of happening, so you should be 87.5% likely to win. But you’re not. Remember that the first 2 losses only has a 25% chance to happen in the first place. 12.5% is, not surprisingly, exactly half of 25%, keeping your current odds squarely at 50-50.

Therefore, once you lose 2 in a row, you have painted yourself into a 25% total probability zone: 12.5% of the time you win, 12.5% of the time you lose, and 75% of the time you never got here in the first place.

I hope that makes sense to you. This was the only way I could shake that pesky gambler’s fallacy.

---

<div class="post-metadata">

**Author:** ![ccwaterback](https://avatars.discourse-cdn.com/v4/letter/c/df705f/32.png) [@ccwaterback](https://boards.straightdope.com/u/ccwaterback)\
**Post date:** [June 30, 2004, 6:18am UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/26 "2004-06-30T06:18:46Z")

</div>

> [@Ellis Dee](#):
>
> Therefore, once you lose 2 in a row, you have painted yourself into a 25% total probability zone: 12.5% of the time you win, 12.5% of the time you lose, and 75% of the time you never got here in the first place.
> 
> I hope that makes sense to you. This was the only way I could shake that pesky gambler’s fallacy.

That’s the reply I was looking for, thanks.

---

<div class="post-metadata">

**Author:** ![TommyTutone](https://avatars.discourse-cdn.com/v4/letter/t/c4cdca/32.png) [@TommyTutone](https://boards.straightdope.com/u/TommyTutone)\
**Post date:** [June 30, 2004, 2:05pm UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/27 "2004-06-30T14:05:58Z")

</div>

> [@Ellis Dee](#):
>
> First, let me state clearly that I know nothing about blackjack or card counting. If your system is predicated on dependent trials, then my logic may not apply. (Though, I suspect, your trials are more independent than you think.)The bolded (by me) part is irrelevant. It also happens to be the core of the gambler’s fallacy.

Unfortunately (or fortunately), in blackjack, events are not independent of one another. If, for example, you lose four times in a row, and the deck has not been shuffled during those four losses, then you have a better chance in winning (in general). That is because you are generally losing when smaller numbers are being played out of the deck, which leaves the larger numbers behind, which is to the player’s advantage (because of the 3-2 payoff on blackjack).

Of course, once the decks are shuffled, everything starts anew and you have to disregard any previous wins or losses.

---

<div class="post-metadata">

**Author:** ![hammos1](https://avatars.discourse-cdn.com/v4/letter/h/ba8739/32.png) [@hammos1](https://boards.straightdope.com/u/hammos1)\
**Post date:** [June 30, 2004, 2:20pm UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/28 "2004-06-30T14:20:27Z")

</div>

> [@ccwaterback](#):
>
> That’s the reply I was looking for, thanks.

I’m genuinely curious, and I hope you don’t mind me asking- **ccwaterback** , your strategy was based on the assumption that, after losing three hands in a row, you had a very low chance of losing a fourth. Can you explain how you thought this to be true? The cards would have to ‘remember’ that you had lost three hands, and arrange themselves so that you’d probably win the fourth…?

---

<div class="post-metadata">

**Author:** ![ccwaterback](https://avatars.discourse-cdn.com/v4/letter/c/df705f/32.png) [@ccwaterback](https://boards.straightdope.com/u/ccwaterback)\
**Post date:** [June 30, 2004, 10:25pm UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/29 "2004-06-30T22:25:07Z")

</div>

> [@hammos1](#):
>
> I’m genuinely curious, and I hope you don’t mind me asking- **ccwaterback** , your strategy was based on the assumption that, after losing three hands in a row, you had a very low chance of losing a fourth. Can you explain how you thought this to be true? The cards would have to ‘remember’ that you had lost three hands, and arrange themselves so that you’d probably win the fourth…?

I think I was just looking at the odds that **Sam** pointed out, at face value.

> [@](#):
>
> But call it 50% for making the math easier. The odds of losing two in a row are then 25%. Three in a row is 17.5%. Four in a row is 8.75%. Etc

But as **Ellis Dee** clarified …

> [@](#):
>
> You just lost 2 in a row, which in itself only has a 25% chance to occur. That means that you already realized the low percentage chance you are trying to capitalize on. As in, you missed your chance.

---

<div class="post-metadata">

**Author:** ![jsc1953](https://avatars.discourse-cdn.com/v4/letter/j/5daacb/32.png) [@jsc1953](https://boards.straightdope.com/u/jsc1953)\
**Post date:** [June 30, 2004, 11:06pm UTC](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683/30 "2004-06-30T23:06:58Z")

</div>

> [@I Love Me](#):
>
> Josh–
> 
> Heck, if I could find a single deck game with a .20 edge for the house I’d love to play it. The only reason I play from a shoe is because those are the games with the best advantage I can find. I’d gladly sacrifice DAS and move to single deck if I knew where the game was. Especially if the penetration was better than half a deck.

The most common occurence of parameters in Reno (for example, at the El Dorado Hotel & Casino) is double-deck, double on anything, no DAS, unlimited re-splits except aces, dealer hits soft 17, 3-2 on blackjack.

[Previous page](https://boards.straightdope.com/t/blackjack-consecutive-wins-or-losses/252683.md?page=1)
